# PDV: Prompt Directional Vectors for Zero-shot Composed Image Retrieval

### v1

@@ 1 Introduction
Figure 1: Figure 1: Overview of Prompt Directional Vector (PDV) for Zero-Shot Composed Image Retrieval (ZS-CIR). (a) Standard ZS-CIR pipeline. (b) PDV calculation process. (c) Dynamic text embedding composition using PDV. (d) Fusion of composed embeddings: PDV-modified image embedding combined with composed text embedding.
@@
ZS-CIR leverages VLMs, denoted by $\Psi$, which operate through a dual-pathway architecture. The first pathway consists of a vision branch, $\Psi_{I}$, that extracts feature embeddings from target images, $I_{target}$. The second pathway employs a language branch, $\Psi_{T}$, that processes a textual composition of reference images, $I_{ref}$, and user-provided text prompts, $P$. This composition, represented by $\mathcal{F}(I_{ref},P)$, can be achieved through two primary methods: (1) Caption Generation, where a caption is generated for the reference image using a VLM, and this caption is merged with the text-prompt using Large Language Models (LLMs), as demonstrated in CIReVL [11]; or (2) Pseudo Tokenization, which uses CLIP’s [16] visual branch to process the $I_{ref}$ and a mapping network (consisting of a lightweight multi-layer perceptron) to tokenise the image, as demonstrated in Pic2Word [18]. The resulting $\mathcal{F}(I_{ref},P)$ is a textual query representation that encompasses both the provided visual and text information, and facilitates zero-shot retrieval. The aforementioned pipeline is illustrated in Figure 1a.
Despite promising results reported in the literature [18, 11, 8, 2], existing approaches include three significant limitations:
Gap 1: Staticly Composed Text Embedding. Target images, $I_{target}$, may not appear in the Top-K retrieved results when other gallery images produce embeddings that are closer to the composed text embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ than the target image embedding $\Psi_{I}(I_{target})$, as shown in Figure 1a. In such cases, users must iteratively refine their prompts and regenerate composed text embeddings, incurring additional manual effort and computational overhead.
Gap 2: Underutilisation of Reference Image Embedding. Current methods typically do not leverage the embedding $\Psi_{I}(I_{ref})$ of the reference image. This omission stems from consistently poor retrieval performance observed when incorporating these embeddings, as documented in multiple studies [18, 11, 2].
@@
Promp Directional Vector (PDV): a Plug-and-Play Solution. We propose a straightforward, training-free approach to address both these challenges by introducing the Prompt Directional Vector (PDV). Denoted by $\Delta_{PDV}$, the PDV represents the residual vector between two text embeddings: the composed text embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ and the reference image text embedding $\Psi_{T}(\mathcal{F}(I_{ref}))$. The latter is equivalent to $\Psi_{T}(\mathcal{F}(I_{ref},P_{Empty}))$, where $P_{Empty}$ represents an empty input string, corresponding to the unprompted baseline. As illustrated in Figure 1b and shown via a red arrow, this PDV captures the semantic modification induced by the prompt. In the following sections, we demonstrate how the PDV effectively addresses the three aforementioned challenges.
@@
PDV: Addressing Gap 1. To change the static nature of the composed text embedding, we generalise the synthesis of the composed text embeddings $\Psi_{T}(\mathcal{F}(I_{ref},P))$. We interpret $\Psi_{T}(\mathcal{F}(I_{ref},P)$ as a shift from the reference image text embedding without the prompt, $\Psi_{T}(\mathcal{F}(I))$, by a vector $\Delta_{PDV}$. Under this formulation, the baseline ZS-CIR approach can be viewed as a special case where $\Psi_{T}(\mathcal{F}(I_{ref},P))=\Psi_{T}(\mathcal{F}(I_{ref},P))+\alpha\Delta_{PDV}$ with $\alpha=1$. We hypothesize that when $\Delta_{PDV}$ captures the desired modifications but not their precise magnitude (particularly with less descriptive prompts), adjusting $\alpha$ can enhance retrieval performance and controllability. As demonstrated in Figure 1c, increasing $\alpha$ to 1.3 produces results more closely aligned with the target compared to the baseline case of $\alpha=1$.
@@
PDV: Addressing Gap 2. Although image embeddings $\Psi_{I}(I_{ref})$ contain valuable visual content regarding the reference image, they lack prompt-specific semantic information, leading to poor performance when used in ZS-CIR. By leveraging the shared semantic space learned by Vision-Language models across image and text modalities, we can transfer prompt semantics to the image embedding by adding the Prompt Vector $\Delta$, obtaining $\Psi_{I}(I_{ref})+\alpha\Delta_{PDV}$, as illustrated in Figure 1d. We denote this augmented representation as the composed image embedding. Similar to the dynamic composed text embedding, this representation can be adjusted through a scaling factor, $\alpha$.
@@
PDV: Addressing Gap 3. Lastly, several studies demonstrate that the direct fusion of image and text embeddings outperforms using either input feautre (image or text-prompt) alone. [18, 11, 2]. However, this fusion approach still underperforms compared to using the composed text embeddings. This performance gap exists because prompt embeddings are significantly changed by incorporating context from the reference image. Specifically, $\Delta$ is not equivalent to $\Psi_{T}(P)$. To address this, we propose fusing the composed text and composed image embeddings, as illustrated in Figure 1d. Through varying the fusion weight factor $\beta$, we can control the balance between visual similarity to the reference image and semantic alignment with the prompt. Lower $\beta$ values prioritize visual fidelity, while higher values emphasize semantic modifications specified in the prompt.
@@
PDV serves as a plug-and-play enhancement for most ZS-CIR approaches, offering a simple and training-free solution. The computational overhead is minimal, requiring only the calculation of text and image embeddings from the reference image. We evaluate PDV by integrating it with two distinct ZS-CIR methods across various CIR benchmarks. Our experimental results demonstrate that all three use cases of PDV consistently improve upon the baseline approaches, particularly when the baseline method already generates accurate compositional embeddings.
@@
  • We introduce the Prompt Directional Vector (PDV), a simple and training-free enhancement for overcoming the limitations of current Zero-Shot Composed Image Retrieval.
  • We propose three novel applications of PDV: (1) dynamic composed text embedding synthesis through PDV scaling, which enhances control over retrieval results, (2) composed image embedding synthesis via semantic transfer from prompts to visual features through PDV addition, which prioritizes visual similarity, and (3) effective fusion of composed text and image embeddings, which improves overall performance and enables controllable balancing of visual and semantic similarity.
  • Through extensive experiments on multiple benchmarks with two ZS-CIR methods, we demonstrate that PDV consistently improves retrieval performance with minimal computational overhead.
@@ 2 Related Work
Vision-Language (VL) models have revolutionized computer vision by effectively bridging visual and textual modalities. The emergence of powerful models such as CLIP [16], ALIGN [10], and Florence [23] has enabled remarkable advances in multi-modal understanding. Trained on large-scale image-text pairs through contrastive learning, these models learn rich visual-semantic representations that generalize across domains and tasks. Building upon these advances, Composed Image Retrieval (CIR) has shown significant progress [4, 18, 11]. Early approaches leverage VL models and either trained a combiner network to compose text and image features [4], or fine-tuned a text encoder [3] to extract task-specific text features. However, these methods still require expensive domain-specific triplets (reference image, modified image, and text description) that must be manually verified. Recent work has explored alternative approaches to reduce the data collection burden, such as using synthetic triplets [9] or mining triplets from large-scale image-text datasets [14]. However, these methods still incur significant computational costs during training.
Zero-shot CIR with Text Inversion Recent research has focused on zero-shot approaches to address these challenges. Many methods adopt text inversion, a technique initially proposed for personalized image generation [7, 17], which maps images to pseudo tokens or words. Pic2Word [18] introduced a self-supervised text inversion network trained with cyclic contrastive loss, though it requires a large-scale image dataset. SEARLE [2] reduces the cost of training Pic2Word and improves the efficiency of the text inversion network. To further improve scalability, LinCIR [8] proposed a language-only approach that reduces training costs and increases scalability. Most recently, CIReVL [11] introduced a more direct method that leverages image captioning models to generate natural language descriptions of reference images, which are then combined with text that specifies desired modifications to form queries.
@@
Comparative Summary. While residual embeddings were widely explored in early CIR studies, they were limited to supervised settings. Recent zero-shot approaches have demonstrated that effective CIR systems can be built by leveraging pre-trained VL models without expensive supervised training. However, key challenges persist in controllability, accuracy, and effective multi-modal fusion, which we address in this work.
3 Improved Zero-shot Composed Image Retrieval (ZS-CIR)
@@ 3.1 Baseline ZS-CIR Framework
Composed Image Retrieval (CIR) enables users to search for target images $I_{target}$ by providing a reference image, $I_{ref}$, and a text prompt, $P$, describing desired modifications. Zero-shot composed image retrieval (ZS-CIR) leverages Vision-Language (VL) models, $\Psi$, such as CLIP [16], whose vision branch, $\Psi_{I}$, and text branch, $\Psi_{T}$, are trained to learn a shared embedding space where semantically similar image and text pairs are mapped close to each other. In this framework, target images are encoded using the vision branch, $\Phi_{I}$, while the query is composed by processing both $I_{ref}$ and $P$ through the text branch $\Psi_{T}$, as composition operations are more naturally handled in the text modality.
Recent ZS-CIR approaches generate the composed text embedding from $I_{ref}$ and $P$ using one of two methods: direct image captioning (CIReVL) or pseudo tokenization (Pic2Word). We denote this composition process as $\mathcal{F}$, resulting in a composed text embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$.
@@
If $I_{target}\notin\mathbb{I}_{top-k}$, the user must reformulate the prompt and repeat the feature extraction process to obtain alternative retrieval results. Notably, as shown in Eq. 1, only the composed feature embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ directly influences the computation of $\mathbb{I}_{top-k}$ results. Although the gallery images are represented by their image embeddings, the image embedding of the reference image $\Psi_{I}(I_{ref})$ does not contribute to the retrieval process.
3.2 Prompt Directional Vector
Rather than simply employing the embedding point alone, $\Psi_{T}(\mathcal{F}(I_{ref},P))$, we propose a generalized formulation of composed text embeddings by considering an embedding direction, $\Delta_{PDV}$, which is derived from the difference between the provided prompt, $P$, and the reference image, $I_{ref}$. Formally, we define $\Delta_{PDV}$ as,
@@ 3.2 Prompt Directional Vector
$\Psi_{T}(\mathcal{F}(I_{ref},P))=\Psi_{T}(\mathcal{F}(I_{ref}))+\alpha\Delta_{PDV},$ (3)
@@
3.3 Strategies for Using the Prompt Directional Vector
We explore three strategies to incorporate the Prompt Directional Vector, $\Delta_{PDV}$.
(1) Prompt Directional Vector for Text (PDV-T), which enhances controllability in ZS-CIR. While baseline ZS-CIR approaches represent a special case where $\alpha=1$, varying $\alpha$ provides users with additional control over the retrieval process. Setting $\alpha>1$ amplifies the modification specified by the prompt, while $\alpha<1$ reduces its effect. This approach offers a more efficient alternative to modifying the prompt directly, as it requires neither new feature extraction nor prompt reformulation. Note we use the notation $\Phi_{PDV-T}$ to represent the composed text embedding.
@@ 3.3 Strategies for Using the Prompt Directional Vector
(2) Prompt Directional Vector for Image (PDV-I), which extends the modification principle to visual embeddings. While previous approaches primarily relied on composed text embeddings, experimental results show that direct fusion of image and text features yields inferior performance compared to composed features. We hypothesize that this performance gap arises because the direct text embedding, $\Phi_{T}(P)$, differs significantly from the prompt vector $\Delta_{PDV}$, as illustrated in Figure 2. This difference occurs because the semantic meaning of natural language is context-sensitive, where in our case the context is provided by the reference image embedding $\Psi_{T}(\mathcal{F}(I_{ref}))$. To address this limitation, we propose incorporating $\Delta_{PDV}$ into visual embeddings. Specifically, we compute the composed visual embedding $\Phi_{PV-T}$ as $\Psi_{I}(I_{ref})+\alpha\Delta_{PDV}$, where $\Psi_{I}(I_{ref})$ represents the original visual embedding of the reference image, and the same prompt vector obtained via Eq. 2 is used to modify this visual representation.
(3) Prompt Directional Vector Fusion (PDV-F), which calculates the final similarity score between a query and target image which combines both composed embeddings. This fusion embedding, $\Phi_{PV-F}$, can be defined as,
$\Phi_{PV-F}=(1-\beta)\Phi_{PV-I}+\beta\Phi_{PV-T},$ (4)
@@ 4.1 Implementation Details
4.2 Datasets and Base Vision-Language Models
Following previous work, we evaluated our method on a suite of datasets including Fashion-IQ [21], CIRR [15] and CIRCO [2]. Our proposed method is a plug-and-play approach requiring no additional training, leveraging only pre-trained models. For feature extraction, we use three CLIP variants: ViT-B/32, ViT-L/14, and ViT-G/14, and used the same pre-trained weights as used by the baseline methods. For image tokenization, we employ the pre-trained Pic2Word model.

### v2

@@ 1 Introduction
Figure 1: Overview of Prompt Directional Vector (PDV) for Zero-Shot Composed Image Retrieval (ZS-CIR). (a) Standard ZS-CIR pipeline. (b) PDV calculation process. (c) Dynamic text embedding composition using PDV. (d) Fusion of composed embeddings: PDV-modified image embedding combined with composed text embedding.
@@
ZS-CIR leverages VLMs, denoted by $\Psi$, which operate through a dual-pathway architecture. The first pathway consists of a vision branch, $\Psi_{I}$, that extracts feature representations from target images, $I_{target}$. The second pathway employs a language branch, $\Psi_{T}$, that processes a textual composition of reference images, $I_{ref}$, and user-provided text prompts, $P$. This composition, represented by $\mathcal{F}(I_{ref},P)$, can be achieved through two primary methods: (1) Caption Generation, where a caption is generated for the reference image using a VLM, and this caption is merged with the text-prompt using Large Language Models (LLMs), as demonstrated in CIReVL [11]; or (2) Pseudo Tokenization, which uses CLIP’s [16] visual branch to process $I_{ref}$ and a mapping network (consisting of a lightweight multi-layer perceptron) to tokenise the image, as demonstrated in Pic2Word [18]. The resulting $\mathcal{F}(I_{ref},P)$ is a textual query representation that encompasses both the provided visual and text information, and facilitates zero-shot retrieval. The aforementioned pipeline is illustrated in Figure 1a.
We identify three major gaps in the literature, despite the promising results [18, 11, 8, 2]:
Gap 1: Staticly Composed Text Embedding. Target images, $I_{target}$, may not appear in the Top-K retrieved results when other gallery images produce embeddings closer to the composed text embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ than the target image embedding $\Psi_{I}(I_{target})$, as shown in Figure 1a. In such cases, users must iteratively refine their prompts and regenerate composed text embeddings, incurring additional manual effort and computational overhead.
Gap 2: Underutilisation of Reference Image Embedding. Current methods generally do not utilize the embedding $\Psi_{I}(I_{ref})$ of the reference image directly for retrieval; instead, $\Psi_{I}(I_{ref})$ is used solely for composition. This omission stems from consistently poor retrieval performance when incorporating these embeddings, as documented in multiple studies [18, 11, 2].
@@
Promp Directional Vector (PDV): a Plug-and-Play Solution. We propose the Prompt Directional Vector (PDV) as a straightforward, training-free approach to address the aforementioned gaps. Denoted by $\Delta_{PDV}$, the PDV represents the residual vector between two text embeddings: the composed text embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ and the reference image text embedding $\Psi_{T}(\mathcal{F}(I_{ref}))$. The latter is equivalent to $\Psi_{T}(\mathcal{F}(I_{ref},P_{Empty}))$, where $P_{Empty}$ represents an empty input string, corresponding to the unprompted baseline. As illustrated in Figure 1b and shown via a red arrow, this PDV captures the semantic modification induced by the prompt. In the following, we summarize how the PDV effectively addresses the three aforementioned challenges.
@@
PDV: Addressing Gap 1. To change the static nature of the composed text embedding and increase flexibility and utility for users, we generalise the synthesis of the composed text embeddings $\Psi_{T}(\mathcal{F}(I_{ref},P))$. We interpret $\Psi_{T}(\mathcal{F}(I_{ref},P)$ as a shift from the reference image text embedding without the prompt, $\Psi_{T}(\mathcal{F}(I))$, by a vector $\Delta_{PDV}$. Under this formulation, the baseline ZS-CIR approach can be viewed as a special case where $\Psi_{T}(\mathcal{F}(I_{ref},P))=\Psi_{T}(\mathcal{F}(I_{ref},P))+\alpha\Delta_{PDV}$ with $\alpha=1$. We hypothesize that when $\Delta_{PDV}$ captures the desired modifications, but not their precise magnitude (particularly with less descriptive prompts), adjusting $\alpha$ can enhance retrieval performance and controllability. As demonstrated in Figure 1c, increasing $\alpha$ to 1.3 produces results more closely aligned with the target compared to the default $\alpha=1$.
@@
PDV: Addressing Gap 2. Although image embeddings $\Psi_{I}(I_{ref})$ contain valuable visual content regarding the reference image, they lack prompt-specific semantic information, leading to poor performance when used in ZS-CIR. By leveraging the shared semantic space learned by Vision-Language models, we can transfer prompt semantics to the image embedding by adding the Prompt Vector $\Delta$, obtaining $\Psi_{I}(I_{ref})+\alpha\Delta_{PDV}$, as illustrated in Figure 1d. We denote this augmented representation as the composed image embedding. Similar to the dynamic composed text embedding, this representation can be adjusted through a scaling factor, $\alpha$ to offer controllability to enhance retrieval.
@@
PDV: Addressing Gap 3. Lastly, several studies demonstrate that the direct fusion of image and text embeddings outperforms using either input feautre (image or text-prompt) alone. [18, 11, 2]. However, this fusion approach still underperforms compared to using the composed text embeddings. This performance gap exists because prompt embeddings are significantly changed by incorporating context from the reference image. Specifically, $\Delta$ is not equivalent to $\Psi_{T}(P)$. To address this, we propose fusing the composed text and composed image embeddings, as illustrated in Figure 1d. Through varying the fusion weight factor $\beta$, we can dynamically control the balance between visual similarity to the reference image and semantic alignment with the prompt without needing to craft new prompts, or modify reference images. Lower $\beta$ values prioritize visual fidelity, while higher values emphasize semantic modifications specified in the prompt.
@@
PDV serves as a plug-and-play enhancement for most ZS-CIR approaches, offering a simple and training-free solution. The computational overhead is minimal, requiring only the calculation of text and image embeddings from the reference image. We evaluate PDV by integrating it with four distinct ZS-CIR methods across various CIR benchmarks. Our experimental results demonstrate that all three use cases of PDV consistently improve upon the baseline approaches, particularly when the baseline method already generates accurate compositional embeddings.
@@
  • We introduce the Prompt Directional Vector (PDV), a simple and training-free enhancement that overcomes limitations of current Zero-Shot CIR methods.
  • We propose three novel applications of PDV: (1) dynamic composed text embedding synthesis through PDV scaling, which offers enhanced control over retrieval results without tedious prompt modification; (2) composed image embedding synthesis via semantic transfer of prompts to visual features through PDV addition, which prioritizes visual similarity; and (3) effective fusion of composed text and image embeddings, which improves overall performance and enables controllable balancing of visual and semantic similarity.
  • Through extensive experiments on multiple benchmarks with four ZS-CIR methods, we demonstrate that PDV consistently improves retrieval performance with minimal computational overhead.
@@ 2 Related Work
Vision-Language (VL) models have revolutionized computer vision by effectively bridging visual and textual modalities. The emergence of powerful models such as CLIP [16], ALIGN [10], and Florence [26] has enabled remarkable advances in multi-modal understanding. Trained on large-scale image-text pairs through contrastive learning, these models learn rich visual-semantic representations that generalize across domains and tasks. Building upon these advances, Composed Image Retrieval (CIR) has shown significant progress [4, 18, 11]. Early approaches leveraged VL models and either trained a combiner network to compose text and image features [4], or fine-tuned a text encoder [3] to extract task-specific text features. However, these methods required expensive domain-specific triplets (reference image, modified image, and text description) that must be manually verified. Recent work has explored alternative approaches to reduce the data collection burden, such as using synthetic triplets [9] or mining triplets from large-scale image-text datasets [14]. However, these methods still incur significant computational costs during training.
Zero-shot CIR with Text Inversion Recent research has focused on zero-shot approaches to address these challenges. Many methods adopt text inversion, a technique initially proposed for personalized image generation [7, 17], which maps images to pseudo-tokens or words. Pic2Word [18] introduced a self-supervised text inversion network trained with cyclic contrastive loss, though it requires a large-scale image dataset. SEARLE [2] reduces the cost of training Pic2Word and improves the efficiency of the text inversion network. KEDs [19] implicitly models the attributes of the reference images by incorporating a database; thus, tokens obtained through inversion include attributes such as color, object number and layout. To further improve scalability, LinCIR [8] proposed a language-only approach that reduces training costs and increases scalability. Most recently, CIReVL [11] introduced a more direct approach that leverages image captioning models to generate natural language descriptions of reference images, which are then combined with text that specifies desired modifications to form queries. Subsequently proposed methods, such as LDRE [24] and SEIZE [23], leverage multiple captions over a single caption to increase the diversity and also take the semantic increment during the composition into consideration.
@@
3 Methodology
@@ 3.1 Baseline ZS-CIR Framework
Composed Image Retrieval (CIR) enables users to search for target images $I_{target}$ by providing a reference image, $I_{ref}$, and a text prompt, $P$, describing desired modifications. Zero-shot composed image retrieval (ZS-CIR) leverages Vision-Language (VL) models, $\Psi$, such as CLIP [16], whose vision branch, $\Psi_{I}$, and text branch, $\Psi_{T}$, are trained to learn a shared embedding space where semantically similar image and text pairs are mapped close to each other. In this framework, as show in Figure 1a, target images are encoded using the vision branch, $\Phi_{I}$, while the query is composed by processing both $I_{ref}$ and $P$ through the text branch $\Psi_{T}$, as composition operations are more naturally handled in the text modality.
Recent ZS-CIR approaches generate the composed text embedding from $I_{ref}$ and $P$ using one of two methods: direct image captioning (CIReVL, LDRE and SEIZE) or pseudo tokenization (Pic2Word, LinCIR, SEARLE and KEDs). We denote this composition process as $\mathcal{F}$, resulting in a composed text embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$.
@@
If $I_{target}\notin\mathbb{I}_{top-k}$, the user must reformulate the prompt and repeat the feature extraction process to obtain alternative retrieval results, incurring time and computational resource costs. Notably, as shown in Eq. 1, only the composed feature embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ directly influences the computation of $\mathbb{I}_{top-k}$ results. Although the gallery images are represented by their image embeddings, the image embedding of the reference image $\Psi_{I}(I_{ref})$ does not contribute to the retrieval process.
3.2 Our Approach: Prompt Directional Vector
Rather than simply employing the composed embedding alone, $\Psi_{T}(\mathcal{F}(I_{ref},P))$, as depicted in Figure 1b, we propose a generalized formulation of composed text embeddings by considering the embedding modification direction, $\Delta_{PDV}$, which is derived from the difference between the provided prompt, $P$, and the reference image, $I_{ref}$. Formally, we define $\Delta_{PDV}$ as,
@@ 3.2 Our Approach: Prompt Directional Vector
$\Psi_{T}(\mathcal{F}(I_{ref},P))=\Psi_{T}(\mathcal{F}(I_{ref}))+\alpha_{T}\Delta_{PDV},$ (3)
@@
3.3 Strategies for Using PDV
We explore three strategies for using $\Delta_{PDV}$:
(1) Prompt Directional Vector for Text (PDV-T), which enhances controllability in ZS-CIR. While baseline ZS-CIR approaches represent a special case where $\alpha=1$, varying $\alpha$ provides users with additional control over the retrieval process (refer to Figure 1c). Setting $\alpha>1$ amplifies the modification specified by the prompt, while $\alpha<1$ reduces its effect. This approach offers a more efficient alternative to modifying the prompt directly, as it requires neither new feature extraction nor prompt reformulation. Note that we use the notation $\Phi_{PDV-T}$ to represent the composed text embedding.
@@ 3.3 Strategies for Using PDV
(2) Prompt Directional Vector for Image (PDV-I), which extends the modification principle to visual embeddings. While previous approaches primarily relied on composed text embeddings, experimental results show that direct fusion of image and text features yields inferior performance compared to composed features. We hypothesize that this performance gap arises because the direct text embedding, $\Phi_{T}(P)$, differs significantly from the prompt vector $\Delta_{PDV}$, as illustrated in Figure 2. This difference occurs because the semantic meaning of natural language is context-sensitive, where in our case the context is provided by the reference image embedding $\Psi_{T}(\mathcal{F}(I_{ref}))$. To address this limitation, we propose combining $\Delta_{PDV}$ with visual embeddings. Specifically, we compute the composed visual embedding $\Phi_{PDV-I}$ as $\Psi_{I}(I_{ref})+\alpha_{I}\Delta_{PDV}$, where $\Psi_{I}(I_{ref})$ represents the original visual embedding obtained from the reference image, and the same prompt vector obtained via Eq. 2 is used to modify this visual representation.
(3) Prompt Directional Vector Fusion (PDV-F), which calculates the final similarity score between a query and target image which combines both composed embeddings. This fusion embedding, $\Phi_{PDV-F}$, can be defined as,
$\Phi_{PDV-F}=(1-\beta)\Phi_{PDV-I}+\beta\Phi_{PDV-T},$ (4)
@@
4 Experiments
[…]
Datasets and Base Vision-Language Models. Following previous work, we evaluated our method on a suite of datasets including Fashion-IQ [22], CIRR [15] and CIRCO [2]. Our proposed method is a plug-and-play approach requiring no additional training, leveraging only pre-trained models. For feature extraction, we use three CLIP variants: ViT-B/32, ViT-L/14, and ViT-G/14, and used the same pre-trained weights as used by the baseline methods. For image tokenization, we employ the pre-trained Pic2Word model.